The preventive advantage of eating fish have been reported in numerous studies. A recent cohort study (Sauvaget C, Nagano J, Allen N, et al. Intake of animal products and stroke mortality in the Hiroshima/Nagasaki Life Span Study. International Journal of Epidemiology. 2003;32:536–543) reported that not eating fish increased the risk of stroke. The table below shows the results of this study: Eating Fish Stroke (Yes) Stroke (No) Total Never 82 1549 1631 Almost Daily 23 779 802 Total 105 2328 2433 A. What was the incidence in the exposed in this study? B. What was the incidence in the unexposed in this study? C. What is the rate difference between the those who never eat fish and those who eat fish almost daily? D. What is the relative risk (RR) for those who never eat fish compare with those who eat fish almost daily? E. Please explain the relative risk.
Answer #1

Eating fish and risk of stroke

Eating Fish Stroke(yes) Stroke(no) Total
Never 82 1549 1631
Almost daily 23 779 802
Total 105 2328 2433

A. What was the incidence in the exposed in this study?

Incidence of exposed = a / a + b

No.of stroke cases (a) = 82

No.of non stroke cases (b) = 1549

Incidence of exposed = 82 / 82 +1549

= 82 / 1631

Incidence of exposed  = 0.0503 or 5.03 per 100

B. What was the incidence in the unexposed in this study?

Incidence of unexposed = c / c+d

No of stroke cases (c) = 23

No.of non stroke cases (d) = 779

Incidence of unexposed = c / c +d

= 23 / 23 +779

= 23 / 802

Incidence of unexposed = 0.0287 or 2.87 per 100

C. What is the rate difference between those who never eat fish and those who eat fish almost daily?

Rate difference is calculated by subtracting the incidence of unexposed form the incidence of exposed cases.

Rate difference = Incidence of exposed cases - Incidence of unexposed cases

= 5.03 - 2.87

Rate difference = 2.16

D. What is the relative risk (RR) for those who never eat fish compare with those who eat fish almost daily?

Relative Risk = Incidence in the exposed / Incidence in unexposed

= a / (a +b) / c / (c + d)

= 5.03 / 2.87

Relative risk = 1.75

E. Please explain the relative risk.

The persons who never eat fish have 1.75 times increased risk of stroke than the persons who eat fish almost daily.